OpenStudy (anonymous):

Discrete. Prove or disprove that " a product of any two rational numbers is rational." I think I would use a direct proof....

jimthompson5910 (jim_thompson5910):

hint: a number x is rational if x = a/b where 'a' and 'b' are integers (b is not equal to 0)

OpenStudy (anonymous):

So take 2 rational numbers x and y. Since x and y are rational, we can write each as a ratio of integers. So let x = a/b and y = c/d. If we multiply those two do we still get a divison of 2 integers :)